OpenStudy (anonymous):

*Will give medal!* The volume of a rectangular prism is 240 cubic feet. The length is 8 feet and the height is 5 feet. What is the width? A. 5 ft B. 6 ft C. 8 ft D. 30 ft

OpenStudy (minecrafter64):

8 ft C

OpenStudy (anonymous):

yeah, its 8

OpenStudy (mathlegend):

To find the volume doesn't it go something like... L * W * H = 240 They gave us... two variables 8 * W * 5 = 40W 40w = 240

OpenStudy (minecrafter64):

whait no its 6 ft

OpenStudy (lone001):

240/5*8=6

OpenStudy (anonymous):

Thanks, that seems so easy now.